A 40.0-mL solution contains 0.033 M barium chloride (BaCl2). What is the minimum concentration of sodium sulfate (Na2SO4) required in the solution to produce a barium sulfate (BaSO4) precipitate? The solubility product for barium sulfate is Ksp = 1.1 ✕ 10−10.
Answer #1

BaSO4 (s) ........> Ba^2+ (aq) + SO4^2- (aq)

Ksp = [Ba^2+][SO4^2-]

or

1.1 * 10^-10 = 0.033 * [SO4^2-]

or

[SO4^2-] = 3.33 * 10^-9 M

thus

minimum concentration of sodium sulfate (Na2SO4) required = 3.33 * 10^-9 M